Find the first fundamental form of surface of revolution?


1
Expert's answer
2021-05-24T03:45:22-0400

 The First Fundamental Form

(note : bold symbol denote to vector, example X denote X vector)

Suppose M is a surface determined by X(u, v) ⊂ E3 and suppose α(t) is a curve on M, t ∈ [a, b]. Then we can write α(t) = X(u(t), v(t)) (then (u(t), v(t)) is a curve in R2 whose image under X is α). Then


α'(t) = (∂X/∂u)(du/dt) + (∂X/∂v)(dv/dt) = u' X1 + v' X2


If s(t) represents the arc length along α (with s(a) = 0) then


s(t) = ∫ || α' (r) || dr

where lower limit is a

and upper limit is t


and

(ds/dt) = || α' (r) ||


so


(ds/dt)2 = || α' (r) || 2 = α' . α' = ( u' X1 + v' X2 ).( u' X1 + v' X2 )

= u'2 ( X1 . X1) + 2 u' v' ( X1 . X2) + v'2 ( X 2. X2)


Following Gauss’ notation (briefly) we denote 


E = X1 . X1

F =X1 . X2

G = X 2. X2


and have

(ds/dt)2 =E (du/dt)2 +2F (du/dt)(dv/dt) + G (dv/dt)2


or in differential notation


(ds)2 =E (du)2 +2F (du)(dv) + G (dv)2

it is represent " first fundamental form of surface of revolution "